OpenStudy (yamyam70):

calculus help Get the conditions of the absolute value function f(x) = |x| + |x - 1| I can individually get the conditions of each term, using the rule |x| = x >= 0 and |-x| x < 0 Im having a hard time getting the specific condition of the whole function, help is much appreciated :)

OpenStudy (anonymous):

By the definition of the absolute value, \[|x|=\begin{cases}x&\text{for }x>0\\0&\text{for }x=0\\-x&\text{for }x<0\end{cases}\] and more generally, \[|x-c|=\begin{cases}x-c&\text{for }x>c\\0&\text{for }x=c\\-(x-c)&\text{for }x<c\end{cases}\] Setting \(c>0\), if we add the two, we have 5 cases: \[\color{red}{|x|}+\color{blue}{|x-c|}= \begin{cases}\color{red}x+\color{blue}{x-c}&\text{for }x>c\\ \color{red}c+\color{blue}0&\text{for }x=c\\ \color{red}x+\color{blue}{c-x}&\text{for }0<x<c\\ \color{red}0+\color{blue}{c-x}&\text{for }x=0\\ \color{red}{-x}+\color{blue}{c-x}&\text{for }x<0\end{cases}\] Simplifying a bit, you have \[|x|+|x-c|=\begin{cases}2x-c&\text{for }x>c\\ c&\text{for }x=c\\ c&\text{for }0<x<c\\ -(x-c)&\text{for }x=0\\ -(2x-c)&\text{for }x<0\end{cases}\]

OpenStudy (yamyam70):

Thanks much , sorry for the late acknowledgement :)